OpenStudy (jenniferjuice):

What is the equation in point-slope form of the line passing through (−2, 0) and (2, 8)? y = 2(x + 2) y = 2(x − 2) y = −2(x − 2) y = −2(x + 2)

OpenStudy (anonymous):

point slope form =y-y1=m(x-x1)

OpenStudy (anonymous):

m=slope

OpenStudy (mathmate):

First find slope using slope \[a=\frac{y_2-y_1}{x_2-x_1}\] then the equation of a line with slope "a" through point P1 is \[y-y_1=a(x-x_1)\] simplify as required.

OpenStudy (anonymous):

A. is the answer